# Collaborative Pareto Set Learning in Multiple Multi-Objective Optimization Problems

### v1

@@ I Introduction
Multi-objective Optimization Problems (MOPs) frequently occur in many real-world applications, such as drug discovery [1], software engineering [2], etc. In MOPs, the objectives typically conflict with each other, rendering it impossible to optimize all the objectives simultaneously with a single solution [3]. Instead, there are Pareto optimal solutions, each representing a different trade-off among the objectives. The primary goal in solving MOPs is to find a set of these Pareto optimal solutions, known as the Pareto set. The image of the Pareto set in the objective space is termed the Pareto front.
Over the past several decades, researchers have developed a variety of population-based Evolutionary Multi-objective Optimization (EMO) algorithms [4, 5, 6]. These algorithms are designed to optimize a population of solutions, ultimately obtaining a set of solutions that are well distributed along the Pareto front. However, a key limitation of these population-based approaches is that merely a finite set of approximate Pareto solutions can be obtained [7]. In response to this limitation, a novel model-based approach known as Pareto Set Learning (PSL) has recently been proposed [7, 8, 9]. PSL focuses on training a neural network model to approximate the entire Pareto set of a given MOP. In this approach, preference vectors serve as inputs to the model, which maps them to Pareto solutions. These solutions are then used to calculate their objective values, which are subsequently employed to refine the model via an optimization process that utilizes a loss function such as the aggregation functions used in decomposition-based EMO [5].
However, most current PSL approaches are designed to handle merely a single MOP. When confronted with multiple MOPs, these approaches typically learn the Pareto set of each MOP sequentially and independently, resulting in significant inefficiencies in terms of both time and memory. Furthermore, it is worth noting that there exist latent synergies between different MOPs [10]. Such sequential and independent learning approaches fail to effectively capitalize on the potential common relationships across these problems. This neglect is especially notable in model-based approaches, where there is a high likelihood of shareable common representations across various MOPs.
@@
To further explore the relationships among different MOPs, we delve into the working mechanics of CoPSL through experimental investigations, demonstrating the existence of common representations that can be shared among MOPs. Leveraging these collaboratively learned common representations helps the framework better approximate Pareto sets. To evaluate the efficacy of CoPSL, we compare it with widely-used EMO algorithms and state-of-the-art PSL approaches. The experimental results demonstrate that the proposed CoPSL not only operates with remarkable efficiency but also outperforms the comparative algorithms in terms of Pareto set approximation, albeit by a modest yet consistent margin.
@@
  • We study the working principle of CoPSL through experimental investigations. It reveals that there exist shareable representations within CoPSL, which help the framework better approximate the Pareto set of each MOP.
@@ II-B Pareto Set Learning
Pareto set learning utilizes a neural network on the user-defined preferences to effectively approximate the entire Pareto set [7, 8, 12]. Specifically, a preference vector $\mathbf{p}=(p_{1},\ldots,p_{m})\in\mathbb{R}^{m}_{+}$ (with $\sum_{i=1}^{m}p_{i}=1$) indicates the desired trade-off among objectives. The network takes the preference vectors as inputs, with the network outputting tailored solutions. Typically, an aggregation function is employed as the loss function, enabling the optimization of MOPs through gradient-based training [13]. Such an optimization process ensures that the generated solutions are not only high quality but also align closely with the defined preferences. This approach has been successfully applied across various domains, including deep multi-task learning [14, 15], neural multi-objective combinatorial optimization [12], etc.
@@ III Collaborative Pareto Set Learning
We begin by outlining the basics of the CoPSL model (refer to Fig. 1). The architecture of the CoPSL model is inspired by the concept of hard parameter sharing in MTL. It aims to collaboratively learn the Pareto sets of multiple MOPs with fewer parameters, enabling more efficient training and inference. In the CoPSL model, the parameter set is divided into shared parameters and MOP-specific parameters, as illustrated in Fig. 1. Let $h_{\phi}(\cdot)$ denote the CoPSL model with parameters $\phi=\{\phi_{shr},\{\phi^{i}_{spec}\}_{i=1}^{K}\}$, where $\phi_{shr}$ represents shared parameters and $\phi^{i}_{spec}$ represents specific parameters for the $i$-th MOP ($i=1,\ldots,K$).
@@ 8       Update the parameters: $\phi\leftarrow\phi-\eta\nabla_{\phi}\mathcal{L}_{tot}$
Compute the losses of the solution sets. With the generated solution $\mathbf{x}_{i}(\mathbf{p}^{b})$, we can calculate its corresponding objective vector $\mathbf{F_{i}}\big{(}\mathbf{x}_{i}(\mathbf{p}^{b})\big{)}$. These objective vectors are crucial for evaluating the solutions. By leveraging these vectors, a loss function can be designed to accordingly update the model, thereby refining its ability to more accurately approximate the Pareto set. For that purpose, we explore four alternatives for CoPSL as follows:
@@
CoPSL-TCH: This method utilizes the Tchebycheff aggregation function, which is designed to focus on the worst-case deviation from the ideal point, denoted as $\mathbf{z}^{\ast}_{i}$, for the $i$-th objective vector [20]. Its loss function is formulated as:
@@
Update the CoPSL model. The CoPSL model parameters set $\phi$ enables to be updated based on the total loss $\mathcal{L}_{tot}$ (Line 11). Particularly, the set of MOP-specific parameters $\phi^{i}_{spec}$ for the $i$-th MOP is updated according to the corresponding loss $\mathcal{L}_{i}$. This suggests that $\phi^{i}_{spec}$ is fine-tuned based on its associated specific solution set. The updated formula is given by:
@@ IV Understanding Shared Representations in CoPSL
The findings indicate that while there appears to be little correlation between the two MOPs F1 and F3, employing CoPSL models with one or two shared layers can still improve overall performance. This demonstrates that certain representations of preferences (inputs) can be generalized and shared across multiple MOPs, rather than being learned separately from scratch for each individual MOP. Notably, this occurs even when the Pareto sets (outputs) are irrelevant, revealing the existence of an effective shareable representation space for preferences across diverse MOPs with limited inherent relation. Leveraging shared knowledge across different MOPs is beneficial for better approximating the Pareto sets.
Furthermore, we note that introducing only one shared layer modestly improves the HV metrics for both MOPs. However, increasing the number of shared layers beyond one reveals a ‘seesaw’ pattern in the HV values for the respective MOPs. This can be attributed to two primary reasons. Firstly, the model’s input consists merely of two-dimensional preference vectors. In cases where the dimensionality is low, the preference vectors present simpler shared representations that can be effectively captured by a single shared layer. Although additional shared layers have the potential to learn more complex representations, they tend to diminish the model’s ability to map these representations to Pareto solutions. Secondly, the disparity in Pareto sets among various MOPs generates conflicting gradients. As delineated in Equation (9), this conflict adversely affects the updates of shared parameters $\phi_{shr}$. Consequently, the given weight vectors face inherent limitations in optimizing for both MOPs simultaneously, as previously mentioned.
@@ V Experiments
We evaluate collaborative Pareto set learning on multiple suites of diverse MOPs. The experimental outcomes indicate that CoPSL achieves a significant enhancement in efficiency and a slight but consistent improvement in performance compared to conventional EMO algorithms and advanced PSL approaches.
@@ V-A Experimental Setups
Evaluation Metrics. The evaluation of algorithmic performance focuses on three primary aspects. 1) Approximation quality of the Pareto set. We employ the HV indicator, which provides a comprehensive evaluation of both the diversity and convergence of the solution set. For a nuanced analysis, we further examine the log HV difference (i.e., the logarithm of the difference between the HV value of the true/approximate PF and the learned PF) to evaluate the gap between the true/approximate and the learned Pareto front. 2) Algorithmic efficiency. We measure this in terms of the actual runtime required to process a problem suite, which serves as a practical performance metric. For model-based algorithms, we further delve into the computational time complexity by analyzing Floating-point Operations (FLOPs), which offers a theoretical perspective on the computational effort required. 3) Space complexity of algorithms. When it comes to model-based algorithms, we consider the total number of parameters within the model, denoted as Params, as a measure of the theoretical computational space complexity.
@@ V-B Comparisons with Base Methods
CoPSL exhibits superior efficiency compared to all the baselines and demonstrates slightly better performance in terms of approximation capability overall. To be specific, CoPSL significantly outperforms the comparison algorithms in runtime performance and boasts lower theoretical FLOPs and Params than its model-based counterparts. This underscores the efficiency of CoPSL in concurrently learning multiple MOPs. Furthermore, CoPSL shows improvement over all the baselines in all but one MOP (i.e., F3), where it still maintains a slight edge over PSL. As depicted in Fig. 3, CoPSL tends to converge slightly faster than the traditional single-MOP PSL approach. This indicates that CoPSL can effectively enhance both the approximation of the Pareto set and the robustness of the model. Additionally, it can be observed that EMOs can outpace traditional PSL in runtime when running with small population sizes. CoPSL addresses this limitation, extending the upper bounds of model-based Pareto set learning.
@@ V-C Comparisons with SOTA PSL Methods
It should be highlighted that there appears to be little correlation between the real-world problems in the nature of questions and the shape of the Pareto sets. Nonetheless, CoPSL is still capable of concurrently learning an entire suite of MOPs while also learning shared common representations that enhance overall performance. This underscores the remarkable potential of CoPSL to simultaneously process numerous MOPs and still derive valuable shared representations.

### v3

@@ I Introduction
Multi-objective Optimization Problems (MOPs) frequently occur in many real-world applications, such as drug discovery [1], software engineering [2], etc. In MOPs, the objectives are typically in conflict, rendering it impossible to optimize all of them simultaneously with a single solution [3]. Instead, there are Pareto optimal solutions, each representing a different trade-off among the objectives. The primary goal in solving MOPs is to find a set of these Pareto optimal solutions, known as the Pareto set. The image of the Pareto set in the objective space is termed the Pareto front.
Over the past several decades, researchers have developed various population-based Evolutionary Multi-objective Optimization (EMO) algorithms [4, 5, 6]. These algorithms are designed to optimize a population of solutions, ultimately obtaining a well-distributed set of solutions along the Pareto front. However, a key limitation of these population-based approaches is that merely a finite set of approximate Pareto solutions can be obtained [7]. In response to this limitation, a novel model-based approach known as Pareto Set Learning (PSL) has recently been proposed [7, 8, 9]. PSL focuses on training a neural network model to approximate the entire Pareto set of a given MOP. In this approach, preference vectors serve as inputs to the model, which maps them to Pareto solutions. These solutions are then used to calculate their objective values, which are subsequently employed to refine the model via an optimization process that utilizes a loss function, specifically the aggregation functions used in decomposition-based EMO [5].
However, most current PSL approaches are designed to handle merely a single MOP. When confronted with multiple MOPs, these approaches typically learn the Pareto set of each MOP sequentially and independently, resulting in significant inefficiencies in both time and memory. Furthermore, it is worth noting that there exist latent synergies between different MOPs [10]. Such sequential and independent learning approaches fail to capitalize on the potential common relationships across these problems effectively. This neglect is especially notable in model-based approaches, where there is a high likelihood of shareable common representations across various MOPs.
@@
To further explore the relationships among different MOPs, we delve into the working mechanics of CoPSL through experimental investigations, demonstrating the existence of common representations that can be shared among MOPs. Leveraging these collaboratively learned common representations helps the framework better approximate Pareto sets. To evaluate the efficacy of CoPSL, we compare it with widely used EMO algorithms and state-of-the-art PSL approaches. The experimental results demonstrate that the proposed CoPSL operates with remarkable efficiency and outperforms the comparative algorithms in terms of Pareto set approximation, albeit by a modest yet consistent margin.
@@
  • We study the working principle of CoPSL through experimental investigations. It reveals that shareable representations exist within CoPSL, which helps the framework better approximate the Pareto set of each MOP.
@@ II-B Pareto Set Learning
Pareto set learning utilizes a neural network on user-defined preferences to approximate the entire Pareto set [7, 8, 12]. Specifically, a preference vector $\mathbf{p}=(p_{1},\ldots,p_{m})\in\mathbb{R}^{m}_{+}$ (with $\sum_{i=1}^{m}p_{i}=1$) indicates the desired trade-off among objectives. The network takes the preference vectors as inputs, with the network outputting tailored solutions. Typically, an aggregation function is employed as the loss function, enabling the optimization of MOPs through gradient-based training [13]. Such an optimization process ensures that the generated solutions are high quality and align closely with the defined preferences. This approach has been successfully applied across various domains, including deep multi-task learning [14, 15], neural multi-objective combinatorial optimization [12], and so on.
@@ III Collaborative Pareto Set Learning
We begin by outlining the basics of the CoPSL model (refer to Fig. 1). The architecture of the CoPSL model is inspired by the concept of hard parameter sharing in MTL. It aims to simultaneously learn the Pareto sets of multiple MOPs with fewer parameters, enabling more efficient training and inference. In the CoPSL model, the parameter set is divided into shared parameters and MOP-specific parameters, as illustrated in Fig. 1. Let $h_{\phi}(\cdot)$ denote the CoPSL model with parameters $\phi=\{\phi_{shr},\{\phi^{i}_{spec}\}_{i=1}^{K}\}$, where $\phi_{shr}$ represents shared parameters and $\phi^{i}_{spec}$ represents specific parameters for the $i$-th MOP ($i=1,\ldots,K$).
@@ 8       Update the parameters: $\phi\leftarrow\phi-\eta\nabla_{\phi}\mathcal{L}_{tot}$
Compute the losses of the solution sets. With the generated solution $\mathbf{x}_{i}(\mathbf{p}^{b})$, we can calculate its corresponding objective vector $\mathbf{F_{i}}\big{(}\mathbf{x}_{i}(\mathbf{p}^{b})\big{)}$. The objective vectors are crucial for evaluating the solutions. By leveraging these vectors, a loss function can be designed to accordingly update the model, thereby refining its ability to approximate the Pareto set more accurately. For that purpose, we explore four alternatives for CoPSL as follows:
@@
CoPSL-TCH: This method utilizes the Tchebycheff aggregation function, which is designed to focus on the worst-case deviation from the ideal point, denoted as $\mathbf{z}^{\ast}_{i}$, for the $i$-th objective vector [20]. Its loss function is formulated as follows:
@@
Update the CoPSL model. The CoPSL model parameters $\phi$ enable to be updated based on the total loss $\mathcal{L}_{tot}$ (Line 11). Particularly, the set of MOP-specific parameters $\phi^{i}_{spec}$ for the $i$-th MOP is updated according to the corresponding loss $\mathcal{L}_{i}$. This suggests that $\phi^{i}_{spec}$ is fine-tuned based on its associated specific solution set. The updated formula is given by:
@@ IV Understanding Shared Representations in CoPSL
The findings indicate that while there appears to be little correlation between the two MOPs F1 and F3, employing CoPSL models with one or two shared layers can still improve overall performance. It demonstrates that certain representations of preferences (inputs) can be generalized and shared across multiple MOPs, rather than being learned separately from scratch for each individual MOP. Notably, this occurs even when the Pareto sets (outputs) are irrelevant, revealing the existence of an effective shareable representation space for preferences across diverse MOPs with limited inherent relation. Leveraging shared knowledge across different MOPs is beneficial for better approximating the Pareto sets.
Furthermore, we note that introducing only one shared layer improves the HV metrics for both MOPs to a certain extent. However, increasing the number of shared layers beyond one reveals a ‘seesaw’ pattern in the HV values for the respective MOPs. This can be attributed to two primary reasons. Firstly, the model’s input consists merely of two-dimensional preference vectors. In cases where the dimensionality is low, the preference vectors present more straightforward shared representations that a single shared layer can effectively capture. Although additional shared layers have the potential to learn more complex representations, they tend to diminish the model’s ability to map these representations to Pareto solutions. Secondly, the disparity in Pareto sets among various MOPs generates conflicting gradients. As delineated in Equation (9), this conflict adversely affects the updates of shared parameters $\phi_{shr}$. Consequently, as previously mentioned, the given weight vectors face inherent limitations in optimizing for both MOPs simultaneously.
@@ V-A Experimental Setups
Evaluation Metrics. The evaluation of algorithmic performance focuses on three primary aspects. 1) Approximation quality of the Pareto set. We employ the HV indicator, which comprehensively evaluates the diversity and convergence of the solution set. For a nuanced analysis, we further examine the log HV difference (i.e., the logarithm of the difference between the HV value of the true/approximate PF and the learned PF) to evaluate the gap between the true/approximate and the learned Pareto front. 2) Algorithmic efficiency. We measure this in terms of the actual runtime required to process a problem suite, which serves as a practical performance metric. For model-based algorithms, we further delve into the computational time complexity by analyzing Floating-point Operations (FLOPs), which offers a theoretical perspective on the computational effort required. 3) Space complexity of algorithms. Regarding model-based algorithms, we consider the total number of parameters within the model, denoted as Params, as a measure of the theoretical computational space complexity.
@@ V-B Comparisons with Base Methods
CoPSL exhibits superior efficiency compared to all the baselines and demonstrates slightly better performance overall in terms of approximation capability. To be specific, CoPSL significantly outperforms the comparison algorithms in runtime performance and boasts lower theoretical FLOPs and Params than its model-based counterparts. This underscores the efficiency of CoPSL in concurrently learning multiple MOPs. Furthermore, CoPSL shows improvement over all the baselines in all but one MOP (i.e., F3), where it still maintains a slight edge over PSL. As depicted in Fig. 3, CoPSL converges slightly faster than the traditional single-MOP PSL approach. This indicates that CoPSL can effectively enhance both the approximation of the Pareto set and the robustness of the model. Additionally, it can be observed that EMOs can outpace traditional PSL in runtime when running with small population sizes. CoPSL addresses this limitation, extending the upper bounds of model-based Pareto set learning.
@@ V-C Comparisons with SOTA PSL Methods
It should be highlighted that there appears to be little correlation between the real-world problems in the nature of questions and the shape of the Pareto sets. Nonetheless, CoPSL is still capable of concurrently learning an entire suite of MOPs while also learning shared common representations that enhance overall performance. This underscores the remarkable potential of CoPSL to process numerous MOPs simultaneously and still derive valuable shared representations.